OpenStudy (anonymous):

How to solve: Int(ArcTan(x^3)/x^4)dx Answer:-ArcTan(x^3)/(3*x^3) + Log(x) - Log(1 + x^6)/6

OpenStudy (anonymous):

\[ I(n)=\int{\arctan x^{n-1}\over x^n}dx\\ =\int x^{-n}\arctan x^{n-1} dx\\ =x^{-n}\int\arctan x^{n-1} dx-\int\left[{d\over dx}(x^{-n})\int\arctan x^{n-1} dx\right]dx \] now, \[ J(n)=\int\arctan x^{n-1}\cdot1dx\\ J(n)=\arctan x^{n-1}\int1dx-\int\left[{d\over dx}(\arctan x^{n-1})\int1dx\right]dx\\ J(n)=x\arctan x^{n-1}-\int\frac{(n-1)x^n}{x^2(x^n+1)}xdx\\ J(n)=x\arctan x^{n-1}-(n-1)\int\frac{x^{n-1}}{x^n+1}dx\\ J(n)=x\arctan x^{n-1}-(n-1)\frac{\ln(x^n+1)}{n}+C \] going back up, \[ I(n)=x^{-n}J(n)+n\int x^{-n-1}J(n)dx\\ I(n)={J(n)\over x^n}+n\int x^{-(n+1)}[x\arctan x^{n-1}-{n-1\over n}\ln(x^n+1)]dx\\ I(n)={J(n)\over x^n}+n\int{\arctan x^{n-1}\over x^n}dx-(n-1)\int{\ln(x^n+1)\over x^{n+1}}dx\\ 2I(n)={J(n)\over x^n}+{n-1\over n}\left[{\ln(x^{-n}+1)}+{\ln(x^n+1)\over x^n}\right] \]

OpenStudy (anonymous):

further simplifying gives \[ I(n)={x\arctan x^{n-1}\over 2x^n}+{n-1\over 2n}\ln(x^{-n}+1) \]

OpenStudy (anonymous):

problem no.1 , why is arctan -ve??

OpenStudy (anonymous):

@rosho could you check for the signs in there?

OpenStudy (anonymous):

\[ I(n)={\arctan x^{n-1}\over 2x^{n-1}}+{n-1\over2n}\left[\ln(1+x^n)-\ln(x^n)\right] \] this is exactly same as the provided answer but with exactly opposite sign!!!!

OpenStudy (anonymous):

oh .. found the correction... the very last step where I substituted J(n) into I(n)